riles413:

A person 5.5 feet tall stands in line of a shadow cast from the top of the house. The shadow hits the top of the person's head, and continues until the shadow ends on the ground 2.8 feet from the person's shoes. The distance along the ground from the tip of the shadow to the house is 16.4 feet. Find the height of the house

sillybilly123:

 seems like: \(\dfrac{H}{16.4 + 2.8} = \dfrac{5.5}{2.8}\)

sillybilly123:

unless:  meaning \(\dfrac{H}{16.4} = \dfrac{5.5}{2.8}\) :)